OpenStudy (anonymous):

help what is the value of x in proportion x+8/5x-2=3/8?

OpenStudy (anonymous):

x=10 ,I think

OpenStudy (anonymous):

@Aamir_ do u think that right?

OpenStudy (anonymous):

What is the value for x in the proportion x+8/5x-2 =3/8 ? Start by multiplying both sides by 5x-2: x + 8 = 3/8(5x - 2) Subtract 8 from both sides: x = 3/8(5x - 2) - 8 Simplify further: x = (15/8)(x) - 35/4

OpenStudy (anonymous):

so what should i put for the answwer?

OpenStudy (anonymous):

no, the answer is ten i was giving you the explanation

OpenStudy (anonymous):

if you divide 15/8 and 35/4 the answer is ten

OpenStudy (anonymous):

what class is this from?

OpenStudy (anonymous):

ummm cnt say because then theyll suspend us

OpenStudy (anonymous):

ok never mind